### v1

# Membership Inference Attacks on Large-Scale Models: A Survey

Abstract
The adoption of the Large Language Model (LLM) has accelerated dramatically since the ChatGPT from OpenAI went online in November 2022. Recent advances in Large Multimodal Models (LMMs), which process diverse data types and enable interaction through various channels, have expanded beyond the text-to-text limitations of early LLMs, attracting significant and concurrent attention from both researchers and industry. While LLMs and LMMs are starting to spread widely, concerns about their privacy risks are increasing as well. Membership Inference Attacks (MIAs), techniques used to determine whether a particular data point was part of a model’s training set, serve as a key metric for assessing the privacy vulnerabilities of machine learning models. Hu et al. show that various machine learning algorithms are vulnerable to MIA. Despite extensive studies on MIAs in traditional models, there remains a lack of systematic surveys addressing their effectiveness and implications in modern large-scale models like LLMs and LMMs. In this paper, we systematically reviewed recent studies of MIA against LLMs and LMMs. We analyzed and categorized each attack based on their methodology and scenario and discussed the limitations in existing research. Additionally, we examine privacy concerns associated with the fine-tuning process. Finally, we provided some suggestions for future research in this direction.
1 Introduction
Aiming to generate human-like text response, the Large Language Models (LLMs) are machine learning models that are based on transformers and trained on massive amounts of data. First introduced by A. Vaswani et al. [1], transformer architecture includes a self-attention mechanism, which allows it to analyze the relation inside the input parallelly. This algorithm shows a significant improvement in efficiency and performance in Natural Language Processing (NLP) compared to early Long Short-Term Memory (LSTM) networks or Recurrent Neural Networks (RNNs) [2] and, therefore, is widely adapted to various Pre-trained Language Models (PLMs) including GPT-2 [3] and BERT [4]. The concept of fine-tuning was also introduced to adapt PLMs—initially trained on general text corpora—to specific tasks by further training them in domain-specific datasets to enhance its performance based on specific tasks [5]. Further research on PLMs found that by expanding the scale of PLM in both size and dataset, the performance of the models shows great improvement [6]. This observation drove the shift from smaller PLMs to massive-scale LLMs capable of generating more fluent and contextually aware text. Benefiting from the deep learning algorithm, attention awareness provided by transformers, and scale effect of PLM, LLM outperformed most of the AI-driven conversational agents. This advantage has led to the widespread adoption of LLMs in multiple fields, including medical services, mail summarization, and text translation.
LMMs are machine learning models that are capable of handling information from multiple modalities [7]. Although LLMs perform well in text generation, the world that human lives in present information across different spectrums, including vision, sound, and other modalities [8]. To overcome this limitation, research and innovation in multimodal models have surged recently, including Gemini Ultra [9] and GPT-4 [10]. While LLMs primarily rely on transformer-based architectures, multimodal models exhibit greater diversity in input-output configurations and employ a wider range of algorithms. The attempt of adoption of multimodal models can be widely found in multiple fields, including robotics, emotion recognition, and video generation.
Membership Inference Attack (MIA) aims to determine if a data point is from the training dataset of the target machine learning model by observing the behavior of the target model when inferring the data point [11]. The basic concept of MIA was first introduced by N. Homer et al. [12] in the context of genomics, where they demonstrated that an adversary could infer whether a specific genome was included in the database of published genomic statistics. By successfully adopting this attack on Convolutional Neural Network (CNN) classifiers, Shokri et al. [11] further proved that MIA is a practical privacy threat in machine learning fields as well. MIA can lead to severe privacy breaches. For example, if an institute has used its members’ data to train a machine learning model and this model is vulnerable to MIA, an attacker could infer whether specific individuals’ data was used for training, potentially leading to privacy violations. Due to its significant consequence, MIA has become a classic approach to testing and measuring the privacy concerns of machine learning models [13]. As a result, a great number of researchers and companies are putting effort into analyzing different attack and defense strategies.
There are several other attacks in machine learning fields similar to MIA, including Attribute Inference Attacks (AIAs), Model Inversion Attacks, and Property Inference Attacks (PIAs). The AIAs focused on extracting the sensitive attributes—such as gender, race, or age of the individual—of specific data that was used in training [14]. Model Inversion Attack aims to reconstruct the original training data by observing the behavior of the target model [15]. PIAs extract the global statistical characteristics of the training datasets [16]. Compared to these attacks, MIA has several advantages in exposing the privacy breach of the machine learning model. Unlike AIAs and PIAs, which mainly focus on specific features or patterns and are contingent upon particular situations, MIAs expose the full data point that contains all the potential sensitive features. Additionally, while Model Inversion Attacks often result in fuzzy reconstructions and face challenges in accurately representing specific data within a class, MIAs provide clear information regarding the presence of the data points in the training set. In this work, we focus on research in the MIA area.
As LLMs and LMMs continue to evolve and gain widespread adoption, addressing their privacy concerns has become increasingly critical. In the LLM area, various research studies have already been conducted that focus on the general privacy risks of LLM [17, 18]. However, these studies provide limited depth and focus in their analysis of MIA. In the field of multimodal models, apart from the work of M.A. Rahman et al. [19], which examines privacy concerns from an educational perspective, and a general review by S.K. Tetarave et al. [20], there remains a lack of in-depth and systematic analysis. Regarding MIA, the study conducted by H. Hu et al. [21] has summarized MIA against various basic machine learning models. While their survey provides a valuable reference for researchers in the area, the study is up to 2022, and there is a lack of study about MIA on more advanced machine learning models. A similar study was published in 2024 by J. Niu et al. [22], which still only summarizes the research attack up to Jan. 2024 and lacks analysis and focus on advanced models. Table 1 summarizes the limitations of existing surveys and highlights the contributions of our research.
To address existing research gaps, this paper covers the following topics:
1. A comprehensive review of recent development of MIA against LLMs and multimodal models. By analyzing the details and categorizing existing attacks, we aim to identify the gaps in current existing research and hope this survey can provide a reference for future studies in the area. To the best of our knowledge, this is the first paper that summarizes MIA on more advanced machine learning models.
2. An analysis of Fine-Tune Database Membership Inference Attack in advanced machine learning models. Given that fine-tuning technology is widely used in deploying LLMs and LMMs, we refine the attack setting for fine-tuning scenarios and review existing MIA research that targets the fine-tuned database’s exposure.
3. Suggestions for future research. Based on the challenges and research gap identified, we provide several suggestions for further exploration in MIA against modern AI models.
Table 1: Summary of Existing Surveys and Our Survey on Large-Scale Models and MIA Attacks. FT: Fine-Tuning, LLM: Large Language Models, LMM: Large Multimodal Models.
Research Large-Scale Model MIA FT Attack Survey Up
LLM LMM In-Depth In-Depth To
[…]
2 Adversarial Models in Membership Inference Attack
2.1 Adversarial knowledge
Black-Box Access This scenario assumes the target model is a “big black box” for the adversary [23]. In this setting, the attacker only has query API to the target model. Therefore, the knowledge is limited to (1) the query result of the data from the target model, (2) the ground truth of the data, and loss and other properties that can be calculated from (1) and (2). Furthermore, it is reasonable to assume that the attacker has (3) some basic limited general information about the target model, like the model type.
White-Box Access This scenario assumes the attacker has full access to the target model [24]. Besides the data in the Black-Box scenario, the attacker can also access (1) the data inside the target model, including the weight, bias, and result from the activation of each neuron; (2) the full train dataset of the target; and (3) the hyperparameters used during the training of the model like learning rate and batch size.
Gray-Box Access This scenario lies between the Black-Box and White-Box settings, assuming the attacker has an extra limited understanding of the target compared to the Black-Box scenario [22]. Besides being able to query the model, the adversary has limited knowledge of the training dataset, including its data structure, and has a portion of data from it.
Different adversarial assumptions give various levels of access to the adversary. Table 2 below is the summarization of each scenario. Black-box access has the most restricted access and is sometimes too strict for the attacker. Due to the misassumption, some research claims using the black-box setting is actually using the gray-box setting. Gray-box access provides a higher level of flexibility but is still suitable for simulating most real-world attacks. For publicly available machine learning models, obtaining partial knowledge of the training data is often feasible. While both the back-box and gray box are able to simulate real-world attacks, the white-box scenario is impractical for it and is suitable for the internal security evaluations for the developer. Since there is no need to infer membership status if an adversary already has access to the entire training dataset.
Table 2: Summarize of attack scenario
Scenario Target model Dataset (Train) Data (query)
General Detail Structure Data Ground truth Output
[…]
2.2 Attack strategies
Target Model-Based Attack In a target model-based attack, the adversary predicts the membership information primarily based on the behavior of the target model. The attack relies on observable outputs, such as prediction confidence, loss values, or correctness of classification. For example, the attack introduced by S. Yeom et al. [25] uses the correctness and loss values from the target classification model for the prediction, and the label-only MIA introduced by C.A. Choquette et al. [26] uses the sensitivity of posterior probabilities given by the target classifier for the attack.
Reference Model-Based Attack The reference model-based attack requires the adversary to develop an additional reference model/shadow model based on its general or detailed understanding of the target model. The attacker infers the membership information based on the behavior of the target model and the reference provided by the reference model. For example, the shadow model MIA introduced by Shokri et al. [11] trains the attack model based on the multiple showdown models that simulate the target classification model.
Compared to reference model-based attacks that require training one or multiple reference models, the target model-based methods require less computational resources and fewer prior investigations about the target model. However, it’s easier for a reference model-based attack to achieve better attack performance since the reference provides more information and “magnifies” the actual difference between the member and non-member data [27, 28]. As a drawback, such improvement heavily depends on the accurate simulation of the target model and the data structure. Additionally, the reference dataset should be largely disjoint from the original dataset. In real-world scenarios, obtaining such an accurate approximation and ensuring a non-overlapping dataset can be challenging, limiting the practicality of this approach [29].
3 Membership Inference Attack against Large Language Model
3.1 Black-Box & Target Model Based
MIN-KPROB [30] The MIN-KPROB MIA is based on the assumption that the response from non-member datapoint is more likely to have the words in the sentence that are less like the words in the rest of the sentence, which are called “outlier,” compared to the member example. An outlier is determined by tokenizing the whole sentence $x$ and calculating the log-likelihood of its specific token $x_{i}$ to the rest of the sentence. To predict the membership identity of a datapoint, it first selects kof tokens that have minimal log-likelihood from the sentence, denote as $x_{MIN-K\%}=(x_{1},x_{2},\cdots,x_{N})$, and calculates the average of the log-likelihood of this set. The k is a pre-designed number. The following equation shows the formula of the result of MIN-KPROB MIA:
$R=\frac{1}{N}\sum_{i=1}^{N}logp(x_{i}|x)$
The MIA model will claim membership identity if the result is higher than a pre-designed threshold. However, its accuracy depends on the pre-designed K and threshold value, which made this method less practical.
Perplexity-Based MIA [31, 32] This MIA method follows the idea that machine learning models tend to have better accuracy when predicting membership identity. It analyzes the average of the perplexity of the whole sentence $x=(x_{1},x_{2},…,x_{N})$, given by:
$exp\left(-\frac{1}{N}\sum_{i=1}^{N}logf_{\theta}(x_{i}|x_{1},x_{2},\cdots,x_{i})\right)$
The MIA model will accept that a data point is a member if the perplexity is lower than a pre-designed threshold, which means the target model is more confident with its prediction. It has several variances including loss based-MIA and zlib entropy based-MIA. However, this model is safer from a high false-positive rate since a well-optimized LLM is able to achieve a high accuracy even for unseen data.
Sensitivity-Based MIA [33, 34, 35] The sensitivity-based MIA model predicts the membership identity of the input data based on the robustness of the target model’s prediction. Specifically, given an input x to be inferred by the target model, the adversary first generates a reference dataset $x^{\prime}=(x_{1}^{\prime},x_{2}^{\prime},\cdots,x_{N}^{\prime})$ by adding perturbation, including randomly deleting words, changing the case of the character, or adding space inside sentences. The target model $f_{target}(x)$ will infer both the original data and the data from the reference dataset, and the prediction is based on the difference in the performance of the target model between the original and the modified data. A threshold is pre-defined and if the result passes the threshold, the attack model predicts it is from the training dataset.
$R=\frac{1}{N}\sum_{i=1}^{N}(f_{t}(x)-f_{t}(x_{i}^{\prime}))$
This attack method assumes the machine learning model is specifically more confident in the prediction of member data and, therefore, becomes more sensitive and yields a larger gap between the original and the reference data. The limitation of this method is that the threshold of the gap needs to be hand-designed, and it’s hard to maintain the consistency of the perturbation added, which could impact the detection result.
MIN-K%++ MIA [36] The MIN-K%++ is based on the assumption that, after tokenizing the sentence, the tokens that have been used for the training of the target model will be assigned with a higher confidence during the response compared to the possibility prediction of other tokens. To implement this assumption, assuming the whole sentence is $x=(x_{1},x_{2},\cdots,x_{t},\cdots,x_{N})$, it compared the possibility of the target token $x_{t}$ to the probability distribution of the whole vocabulary for the next token by following formula:
$MIN-K\%++(x_{1},x_{2},\cdots,x_{t})=\frac{\log p(x_{t}|x_{1\sim(t-1)})-\mu}{\sigma}$
The $\mu$ and $\sigma$ stand for the expectation and standard deviation of the probability distribution of the next token $x_{(t+1)}$’s log possibility. A higher score indicates that compared to the next token, the target model has higher confidence when calculating the current token, which indicates the current token may be presented in the training dataset and the next may not. After calculating the MIN-K%++ score of each token, similar to MIN-KPROB MIA [30], the system calculates the average of the kof the tokens that have the minimum score as the final score for the sentence. Depending on the pre-designed threshold, the identity could be predicted.
While the MIN-K%++ MIA uses the probability distribution of the next token to calibrate its score, it faces several limitations. Firstly, higher final scores indicate the consistency of the confidence of each token, which could indicate most of the tokens from the training dataset, so the sentence is used in training. However, if the sentence has most tokens not from the training dataset, the final score of that sentence could also be high.
Furthermore, a sentence containing most words presented in the training dataset doesn’t necessarily mean the sentence is in the training dataset.
3.2 Gray-Box & Reference Model Based
Likelihood Ratio MIA [28, 37] The strategy that supports likelihood ratio MIA is to use the likelihood ratio test to distinguish the scenario in which the target data appeared in the training dataset and the scenario in which the target data is independent. The likelihood of the former scenario can be calculated by querying the target model, which parameterized by $\theta$. For the latter scenario, a reference model trained on random data, which parameterized by ${\theta}_{r}$, is used to analyze the likelihood. The likelihood ratio test is implemented by the following equation, where s is the target data.
$\mathit{\Lambda}\left(s\right)=\frac{P\left(\left.s\right|\theta\right)}{P\left(\left.s\right|{\theta}_{r}\right)}$
The prediction is based on comparing $\mathit{\Lambda}\left(s\right)$ to the threshold, which was generated by calculating the likelihood ratio score on general data with around 10% false positive rate tolerance.
Self-calibrated Probabilistic Variation MIA (SPV-MIA) [38] Similar to other referenced model-based MIA, SPV-MIA uses referenced models to calibrate the accuracy of the prediction. However, unlike most other referenced MIA that require a careful pre-investigation to collect the data, which should have a similar distribution but most likely exclusive from the training dataset, for the training of the reference model, SPV-MIA assumes that the response of the target model would have similar data structure to the training dataset and therefore use the generated data from the target LLM to train the reference model. SPV-MIA introduces the probabilistic variation ${\tilde{p}}_{\theta}\left(x\right)$ for the comparison of the target data under the target model and reference model, which can be calculated by following formula:
${\tilde{p}}_{\theta}\left(x\right)\approx\frac{1}{2N}\sum^{N}_{n}{\left(p_{\theta}\left(x+z_{n}\right)+p_{\theta}\left(x-z_{n}\right)\right)-p_{\theta}\left(x\right)}$
The $p_{\theta}\left(x\right)$ is probability of the token $x$ and $x\pm z_{n}$ is the symmetrical text pair of $x$ in different directions. The final result can be predicted by comparing the difference of probabilistic variation with a pre-designed threshold.
The SPV-MIA approach reduced the effort needed to collect reference datasets, and the selection of probabilistic variation metric improves the accuracy of the model. However, the assumption that the target model always generates data similar to the training dataset may have several limitations. The data generated by the target model may have too many features from the training dataset besides data structure, making the reference dataset not random enough. It might also face challenges if the training dataset content multi-structure data.
Data Inference MIA [33] Unlike most MIA that use a single metric to judge the membership identity of the data, Data Inference MIA uses a collection of MIA methods for prediction. To train the attack model, the adversary has to collect a set of suspected data $D_{sus}$ and a set of data that is clearly not used during the training $D_{out}$. It is worth noticing that since $D_{out}$ can be easily generated, it didn’t require the collection of the member data. After collecting the data, the $D_{sus}$ and $D_{out}$ is further separate to $T_{sus}$, $V_{sus}$, $T_{out}$, and $V_{out}$. Using the MIA model including Min-kProb [30], Perplexity Based MIA [31], and several reference-based MIA, a feature map is drawn for each data. Using the data from $T_{sus}$ and $T_{out}$, an attack model is trained to distinguish the data from $D_{sus}$ and $D_{out}$. Then, the attacker uses $V_{sus}$ and $V_{out}$ to certify that this attack model is sufficient and the gap between $D_{sus}$ and $D_{out}$ do exist, which indicates the membership identity of $D_{sus}$.
Data Inference MIA eliminated the necessity of the pre-designed threshold by analyzing the existence of difference between the suspecter and outlier. However, depending on the data structure of the outlier collected, the system might be confused by the irrelevant differences and therefore have a high false-positive rate.
3.3 Gray-Box & Target model based
Semantic MIA [39] Following the sensitivity-based MIA, sematic MIA further controls the content of the perturbation added to input data to improve the attack performance. The attack is based on the similar assumption that the target model has different sensitivity for member and non-member data. However, to make sure the difference of the perturbation added each time does not disturb the attack, besides only considering the difference in prediction results between the original data and noisy data, the semantic distance brought by the perturbation, denote by $\phi\left(x\right)-\phi\left(x^{\prime}\right)$, is also analyzed. A neural network is used as the attack model $f_{attack\_NN}\left(x\right)$ and trained with supervised learning, which requires gray box access to get a portion of the training dataset.
$p=\frac{1}{N}\sum^{N}_{i=1}{f_{attack\_NN}\left(f_{target}\left(x\right)-f_{t}\left(x^{\prime}_{i}\right),\phi\left(x\right)-\phi\left(x^{\prime}_{i}\right)\right)}$
Compared to sensitivity-based MIA, semantic MIA has dynamic adjustment over the different amounts of perturbation added. This approach approves the accuracy and robustness of the original method. The introduction of a neural network for the attack model enhances the capability of the attack model but brings the requirement of less restricted access.
Confidence-Based MIA [40] This Confidence-Based MIA utilizes the principle that target models usually have higher confidence when predicting data that it has seen. In LLM, this concept can be implemented by checking the confidence of the predicted current token given the preview tokens. Firstly, the MIA system attracts and normalizes the probability of each token from the input. Based on the calculated result, a feature map is generated for the input. The feature map will be fed into a classification model, which is trained by the data drawn from the training dataset and the independent data, to predict its membership information.
This MIA approach is less complex and requires less computation resources compared to some advanced algorithms. However, a huge difference in the confidence of member and non-member input is not guaranteed, which could lead to lower accuracy.
Data Inference-Based MIA [41] This MIA system is based on the data inference MIA [33]. Similarly, it deploys multiple MIA models to generate the feature map of each data collected, including perplexity-based MIA [31] and MIN-K%++ [36]. However, instead of training the attack model in the suspected dataset and outlier dataset, this system directly trains the attack model on pre-known member and non-member datasets, which was possible under the gray-box scenario.
Compared to the original attack method, the data inference-based MIA approach reduces the possibility that attack models are disturbed by irrelevant differences between the suspect dataset and outlier dataset and, therefore, have better performance. The drawback is requesting a higher level of access. The idea of embedding a variety of MIA methods makes these kinds of approaches more convincible compared to the others but may provide less contribution in the research dimension.
3.4 White-Box & Target Model Based
Noise Neighbor MIA [42] The Noise Neighbor model predicts the membership information based on the sensitivity-based MIA approach [33]. Given the target input and several other inputs that are similar to the target with a specific distance, which is called noise neighbor in this system, the model will behave more sensitively if the target input is from the training dataset. Therefore, a deep gap can be found between the accuracy of the prediction of the target input and its noise neighbors. In the implementation of noise neighbor MIA, gaussian noise was added to the input after passing the embedded layer, which requires white box access. The identity could be predicted by comparing the perplexity between the target and its neighbors.
Compared to the origin sensitivity-based MIA and variance semantic MIA, noise neighbor achieves an even higher consistency in the perturbation added to the target data. The performance of the noise neighbors MIA model is comparable with the shadow model-based model but requires much fewer computation resources. However, the requirement of accessing the embedded layer may significantly limit its application.
Layer Activation-Based MIA The layer activation-based MIA systems predict membership information by examining the activation’s output of each neural. The LUMIA [43] implement this concept by adding Multi-Layer Perceptron (MLP), called Linear Probes (LP), to the output of each layer. LP generates its prediction in each layer, and LUMIA selects the highest score as final result.
The concept and implementation of the LUMIA are relatively simple and suitable for both LLMs and multimodal models. However, whether Multi-Layer Perceptron on the output of activations is sufficient enough to predict the membership information still needs further investigation. Only selecting the highest score may also make the system more vulnerable to bias or less accuracy.
3.5 Visualization of Taxonomy
To provide readers with a clear view and facilitate rapid navigation of the MIA algorithms being analyzed, Figure 1 presents a visualized categorization based on both attack scenarios and methodology.
[…]
my tree style/.style= for tree=grow=east, parent anchor=east, child anchor=west, where level=0my node=black,text width=7.5em,rotate=90,align=center,anchor=center, where level=1my node=black,text width=6em, where level=2my node=black,text width=6em, where level=3my node=black,text width=15em, l sep=1.5em, forked edge, fork sep=1em, edge=draw=black!50, thick, if n children=3for children= if n=2calign with current , tier/.option=level, {forest} my tree style [MIA against LLM [Methodology [Reference Model[Likelihood Ratio MIA[28, 37]
SPV-MIA[38]
Data Inference MIA[33]]] [Target Model[MIN-K%PROB[30]
Perplexity-Based MIA[31, 32]
Sensitivity-Based MIA[33, 34, 35]
MIN-K%++ MIA[36]
Semantic MIA[39]
Confidence-Based MIA[40]
Data Inference-Based MIA[41]
Layer Activation-Based MIA[43]
Noise Neighbor MIA[42]]] ] [Scenario [White-Box[Layer Activation-Based MIA[43]
Noise Neighbor MIA[42]]] [Gray-Box[Likelihood Ratio MIA[28, 37]
SPV-MIA[38]
Data Inference MIA[33]
Semantic MIA[39]
Confidence-Based MIA[40]
Data Inference-Based MIA[41]]] [Black-Box[MIN-K%PROB[30]
Perplexity-Based MIA[31, 32]
Sensitivity-Based MIA[33, 34, 35]
MIN-K%++ MIA[36]]] ] ]
Figure 1: Visualized Taxonomies for LLM focus MIA
4 Membership Inference Attack against Large Multimodal Model
Unlike LLMs, which primarily encompass transformer-based NLP models, LMMs exhibit greater variability depending on their ability to process different input modalities. Therefore, in surveying MIA against multimodal models, it is essential to consider not only the attack scenarios and methodologies but also the input modalities. Given that input modality plays a more crucial role in determining a model’s capabilities, this survey adopts it as the primary taxonomic criterion.
4.1 Vision-language Model (VLM)
Feature-Based MIA The M4I MIA introduced by P. Hu et al. [44] is a feature-based MIA that explores the relation between the features of the images and texts pair, which is the input and output, for the prediction. Firstly, the attacker develops a neural model, including an image decoder for the feature extraction of the images, a text decoder for the analysis of the text input, and several fully connected layers for the fusion of encoded data, analysis of the relation between the features, and final classification. After that, under the gray box setting, the adversary trains the model with the member and non-member image-text pair. Finally, the attacker provides the attack model with the target input and the target model’s output to infer the identity.
For the limitation of this attack, the assumption that unique distinguishable feature exists between the image and text pair of the training dataset is not guaranteed.
Rényi entropy-based MIA [45] The MaxRényi MIA system utilizes the Rényi entropy to detect membership information. Similar to the other research, it assumes the target model has less perplexity when inferring data from the training dataset, which can be measured by less Rényi entropy. For the image MIA, the adversary first queries the target model with an image and promotion, with the description of the image returned. After that, the attacker queries the target with the image, promotion, and original response. The output logits from the target model are separate to image, instruction, and description logits. The final prediction can be made by calculating the Rényi entropy over the top Klogits. The pipeline introduced utilizes the instruction and description token instead of the image token, which makes it suitable for the attack in VLM since VLM has no image token. For the text MIA, since VLM do have text token, it can be achieved by computing the Rényi entropy over the top Kpositions.
Cosine Similarity-Based MIA [46] This attack algorithm is based on the assumption that the target model enhances the cosine similarity during the training. Therefore, the adversary can predict the identity based on the cosine similarity between the target input $x$ and its output $f_{target}\left(x\right)$, given by:
$CS\left(x,f_{target}\left(x\right)\right)>\tau.$
In the research, two advanced MIAs that are based on cosine similarity are also introduced. Augmentation-Enhanced Attack (AEA) is based on the finding that the cosine similarity of the membership data is more sensitive after applying the data augmentation. Therefore, AEA utilizes this magnifying effect and adds K transformations to the $x$, denote as $T_{k}\left(x\right)$. The final prediction is based on:
$CS\left(x,f_{target}\left(x\right)\right)-CS\left(T_{k}\left(x\right),f_{target}\left(x\right)\right)>\tau$
Unlike the original attack and AEA, which use cosine similarity directly for the prediction, Weakly Supervised Attack (WSA) uses it to collect meta-member datasets. The attacker first queries the target model with a dataset that surely is not used in training, $D_{out}$. The cosine similarity of the data and its answer from the $D_{out}$ follows the Gaussian distribution, allowing the attacker to decide a threshold at the higher end ($\mu+3\sigma$ for example) of the distribution. Such threshold is then applied to the general internet databases that might be used for the training to form a suspect member database $D_{sus}$. Finally, a binary classifier is trained to distinguish the feature of the image-text pair from $D_{out}$ and $D_{sus}$ as the attack model.
For the performance of the three approaches, AEA is marginally better than the base method, while WSA is generally better than AEA. However, WSA requires more computation resources and the knowledge of general databases that might be used for training. Furthermore, the assumption of maximizing cosine similarity may not always be valid depending on the specific algorithm of the target model.
Layer Activation-Based MIA The LUMIA [43] can also attack VLM with similar method mentioned in section 3.4.
Sensitivity-Based MIA The sensitivity-based MIA introduced by J. Ren et al. [34] in VLM is very similar to the one in LLM, which is based on the principle that member data is more sensitive and drops more when perturbation is added. Specifically, it paraphrases the text response as the perturbation and measures the average log-likelihood to justify the fluctuation. To design the threshold, it uses a part of the training dataset under the gray-box access and a reference dataset that was surely not used during the training to calculate the trend of the change of members and non-members at sample size.
Temperature-Based MIA The temperature-based MIA against VLM introduced by Y. Hu et al. [47] is based on the perplexity-based MIA that utilizes the accuracy of the target model for prediction. In their research, they further use temperature to magnify the difference in accuracy between member and non-member data. The research adapts this attack to multiple scenarios, including the situation only target data available, which compares the difference in the accuracy between the high temperature and low temperature with a pre-designed threshold, and the gray-box-like scenario by training the attack model based on the patterns of shadow-model’s members and non-members’ accuracy in different temperatures. Their research also contains a gray-box MIA Reference Inference MIA by comparing the target data with the reference set using the statistical hypothesis test (z-test), and a black-box Consistency-Based MIA that continues to query the target model and checks the consistency of the result with members tending to have more consistent outputs.
The research has analyzed multiple situations with different levels of understanding about the database, which increases the capability of the attack system. However, access to the temperature setting of the target model may already require white-box access since it is normally located in the SoftMax layer. The requirement may let the system less practical.
4.2 Multimodal Emotion Recognition (MER) Model
Representations-Based MIA [48] his white-box-based MIA model utilizes the representations, which are the encoded data of the sub-unimodal that analyzes a specific modality, of the input in each modality for the attack. A dense neural network binary classifier is used as the attack model. The adversary first collects a portion of the training dataset and a certain number of non-members, forming the attack dataset. The attacker fed the attack dataset to the target model to extract the representations of each data. These representations with the labels of member and non-member are used for the training of the attack model.
The attack algorithm is relatively easy to understand and requires less computational resources. However, without specification, a single neural network classifier might have limitations in fully finding the difference between member and non-member.
[…]
my tree style/.style= for tree=grow=east, parent anchor=east, child anchor=west, where level=0my node=black,text width=14em,rotate=90,align=center,anchor=center, where level=1my node=black,text width=6em, where level=2my node=black,text width=6em, where level=3my node=black,text width=15em, l sep=1.5em, forked edge, fork sep=1em, edge=draw=black!50, thick, if n children=3for children= if n=2calign with current , tier/.option=level, {forest} my tree style [MIA against Multimodal Model [Methodology [Reference Model[Temperature-Based MIA[47]]] [Target Model[Feature-Based MIA[44]
Rényi entropy-based MIA [45]Cosine Similarity-Based MIA[46]
Layer Activation-Based MIA[43]
Sensitivity-Based MIA[34]
Temperature-Based MIA[47]
Reference Inference MIA[47]
Consistency-Based MIA[47]
Representations-Based MIA[48]]] ] [Scenario [White-Box[Layer Activation-Based MIA[43]
Temperature-Based MIA[47]
Representations-Based MIA[48]]] [Gray-Box[Feature-Based MIA[44]
Cosine Similarity-Based MIA[46]
Sensitivity-Based MIA[34]
Reference Inference MIA[47]]] [Black-Box[Rényi entropy-based MIA[45]
Cosine Similarity-Based MIA[46]
Consistency-Based MIA[47]]] ] [ Modality [MER[Representations-Based MIA[48]]] [VLM[Feature-Based MIA[44]
Rényi entropy-based MIA [45]
Cosine Similarity-Based MIA[46]
Layer Activation-Based MIA[43]
Sensitivity-Based MIA[34]
Temperature-Based MIA[47]
Reference Inference MIA[47]
Consistency-Based MIA[47]]] ] ]
Figure 2: Visualized Taxonomies for Large Multimodal Model focus MIA
4.3 Visualization of Taxonomy
A visual categorization of the analyzed MIA algorithms, organized by target model’s modality, attack scenario and methodology, is presented in Figure 2 to foster reader comprehension and efficient navigation.
5 Fine-tuning MIA in Advanced Machine Learning
The escalating scale and complexity of advanced machine learning models have propelled the adoption of fine-tuning techniques [41, 49], enabling the efficient deployment of these models across diverse domains and tasks. Unlike traditional methods that involve training models from scratch, fine-tuning leverages knowledge acquired during the pre-training phase, thereby achieving comparable performance with substantially reduced computational resources [50].
In the typical development pipeline, pre-training is normally conducted by model developers, while end-users or specific entities perform fine-tuning to tailor models to their unique applications. This pipeline lets fine-tuning often involves utilizing highly private or user-specific data. MIA aims to expose the data used during the training process. Its adoption in the fine-tuning process introduces an even heightened privacy concern.
In this section, we first refine the attack scenarios to account for the more restricted information accessibility during the fine-tuning stage. Following this, we review existing research efforts aimed at adapting MIAs to retrieve the fine-tuning datasets of LLMs and multimodal models.
5.1 Attack scenario
In the context of fine-tuning, our revised attack scenario definitions largely retain the conventional distinctions of black-box and white-box access, as outlined below:
Black-Box Access The adversary can only query the fine-tuned model, obtaining its inference results. Additionally, the attacker possesses general knowledge about the target model but lacks access to its internal parameters or training data.
White-Box Access The adversary has full access to both the pre-trained and fine-tuned models at the developer level. This includes complete knowledge of the model architecture, parameters, and both the pre-training and fine-tuning datasets.
The definition and characterization of gray-box access require refinement, as fine-tuning in real-world scenarios typically exhibits the following characteristics:
  1. 1.
    Availability of Pre-Trained Models: Many companies, such as OpenAI and Google, provide access to pre-trained models along with fine-tuning APIs for public use. As a result, it is relatively feasible for an attacker to have limited access to the pre-trained model and utilize the fine-tuning API to follow the fine-tune process of the target model.
  2. 2.
    Unavailability of Fine-Tuning Datasets: Since fine-tuning is usually performed by end users or organizations, the corresponding datasets are often well-protected. Consequently, attackers face greater challenges in understanding the structure of the fine-tuning dataset or obtaining portions of its data.
Given these distinctions, we further classify gray-box access into two categories: practical gray-box access and full gray-box access, defined as follows:
Practical Gray-Box Access In this scenario, the adversary has conventional gray-box access to the pre-trained model. This includes knowledge of the training dataset’s structure and partial access to its data. However, for the fine-tuned model, the fine-tuning dataset remains inaccessible due to security measures. The attacker can also utilize the fine-tuning API to replicate a similar fine-tuning process to that of the target model.
Full Gray-Box Access In addition to the privileges available in practical gray-box access, the adversary also possesses gray-box access to the fine-tuned model. This includes knowledge of the fine-tuning dataset structure, partial access to the fine-tuning data, and an understanding of the specific fine-tuning techniques employed in the target model.
Among the defined access settings, black-box and practical gray-box scenarios are considered the more suitable for real-world attack scenarios. In contrast, full gray-box and white-box access provide extensive information that is unlikely to be available to external attackers, making them more relevant for internal security evaluations and research on potential privacy vulnerabilities. Table 3 shows the summarization of the information accessible in each scenario.
Table 3: Summarize of attack scenario
Scenario Model Dataset Data (query)
General Detail
(PT/FT)
API
Structure
Data
Structure
Data
Truth Output
[…]
  • 1
[…]
5.2 Fine-Tuning MIA in Advanced Model
Several studies on MIA against LLMs [42, 35, 32] and multimodal models [43, 34] have included fine-tuned models as their attack targets, demonstrating that MIA techniques can be adapted to fine-tuned settings. However, these studies do not focus on investigating the privacy risks specifically introduced by the fine-tuning process, resulting in limited analysis and a lack of deeper insights.
The first in-depth analysis of fine-tuning’s impact on MIA against LLMs was conducted by F. Mireshghallah et al. [37] in 2022. Using a likelihood ratio-based MIA, they evaluated the vulnerability of fine-tuned LLMs under different fine-tuning strategies, including full fine-tuning, head fine-tuning, and adapter fine-tuning. Their findings indicate that both the extent of fine-tuning and the location of trainable parameters within the model significantly influence MIA susceptibility. Specifically, models fine-tuned on a larger subset of parameters and those where trainable parameters are located closer to the model’s output layer (head) tend to exhibit greater privacy leakage.
After that, the Self-Prompt Calibration MIA (SPV-MIA) introduced by W. Fu et al. [38] uses self-generate content to replace the unpractical a portion of the fine-tune dataset. The study also analyzed the MIA system against LLMs fine-tuned by various Parameter Efficient Fine-Tuning (PEFT) techniques. Their findings also support the observation that MIA vulnerability increases as the number of trainable parameters expands during fine-tuning.
In another study, H. Puerto et al. [41] examined data inference-based MIA at a scaled-up level, targeting continuous learning fine-tuning and end-task fine-tuning in LLMs. Their results indicate that end-task fine-tuning tends to cause greater privacy breaches, particularly when under the collection or database scale level MIA.
For the LMMs, there is still a lack of research about MIA against the fine-tuned model.
One notable study conducted by Z. Li et al. [45] successfully targeted the pre-training dataset in fine-tuned models. Their findings indicate that the MIA vulnerability of the pre-training dataset increases with the number of trainable parameters involved in fine-tuning. Additionally, their results demonstrate that token-based MIAs generally outperform other approaches in detecting membership in fine-tuned models.
Another in-depth study was conducted by Y. Hu et al. [47], which adapted the temperature-based MIA to fine-tuned versions of LLaVA and MiniGPT-4. Their results show that fine-tuned LLaVA exhibited worse MIA resilience compared to MiniGPT-4 under the same conditions. This observation also suggests that the scale of the fine-tuning process may similarly impact privacy vulnerabilities in multimodal models, mirroring the trends observed in LLMs.
The studies analyzed have utilized various settings based on their designed scenario and attack methodology. Table 4 shows the summarization of each research.
Table 4: Summarize of Fine-Tune MIA
Attack LLM/LMM Targets FT Method Scenario Time In-depth
[41] LLM Pythia 2.8/6.9B LoRA Full Gray 2025 Y
[37] LLM GPT-2 Full FT, Adapters, Partial FT Practical Gray 2022 Y
[38] LLM GPT-2/J, Falcon-7B, LLaMA-7B LoRA, Prefix Tuning, P-Tuning Practical Gray 2023 Y
[42] LLM GPT-2 small White 2024 N
[43] LMM LLaVa-OneVision White 2024 N
[45] LMM miniGPT-4, LLaVA, LLaMA Adapter Partial FT, Full FT, Parameter-Efficient FT Black 2024 Y
[34] LMM Pythia-1.4B, LLaVA Full Gray 2025 N
[47] LMM LLaVA, MiniGPT-4 Partial FT Black, Full Gray, White 2025 Y
[35] LLM GPT-2 Black 2023 N
[32] LLM BERT Black 2021 N
  • 1
    –: Not Specified
6 Suggestion and Direction for Future Research
Threshold-Free Black-Box Attack A key limitation observed in current black-box MIA approaches is their reliance on predefined thresholds, which can significantly bias attack performance. In real-world scenarios, accurately estimating such a threshold is either impractical or requires access to a large amount of target data to extract patterns. While the black-box setting is designed to simulate restricted access, the introduction of a threshold reduces the attack’s practicality. Future research should focus on developing threshold-free black-box MIA techniques to improve real-world applicability. Alternatively, investigating whether thresholds derived in controlled experiments can be adapted or transformed for real-world attacks could also be valuable.
Bias From the Database and Model Across the MIA results from different research, the performance of the MIA systems has been found to be heavily influenced by the chosen dataset and target model [29]. Furthermore, most existing studies rely on lab-collected datasets, which are pre-processed, well-balanced, and noise-free. However, it differs significantly from real-world datasets, which tend to be noisier and less balanced. To improve the robustness of MIA models, future research should conduct evaluations across a broader range of datasets and models, including those with imbalanced or noisy data. Furthermore, analyzing how real-world dataset properties could affect the model’s vulnerability to MIA would also provide insights into building more resilient attack and defense strategies.
Real Privacy in MIA Most MIA studies evaluate privacy leakage by measuring the general exposure of training data. However, not all training data points are equally sensitive or privacy-critical. While protection techniques like Differential Privacy (DP) can improve the resilience of the model, they often come with high computational costs and performance degradation. Future research could focus on identifying which parts of the training data are more privacy-sensitive or higher leaked and exploring selective privacy protection mechanisms that target only sensitive data, thereby improving privacy protection while minimizing performance loss.
Cross-Model MIA Most existing MIA techniques are designed for models under specific categories, such as LLMs or VLMs, with only a few exceptions, like LUMIA [43], which can attack both. This may be due to algorithmic differences between models from different categories. However, this limitation reduces the practical utility of MIA systems. Future research should aim to develop MIA methods that utilize features or vulnerabilities across different machine learning models, increasing their general applicability.
MIA Vulnerability Factors in Multimodal Model The LMMs has a higher level of flexibility in terms of algorithms, as they may contain various unimodal algorithms for encoding data from different modalities or deploying different machine learning strategies. Therefore, it would be highly essential and valuable to analyze the impact of the factors on the multimodal models’ MIA vulnerability. These factors include the type of sub-models used for encoding different modalities, the structure and the central algorithm of the model, the distribution of trainable parameters across model components, and other features. Such research will allow the developer to optimize the privacy resilience of the model starting from scratch, promoting more robust model and reducing the reliance on and the defect from post-training privacy mechanisms.
Robust Fine-Tuning MIA While a number of recent studies have started exploring Fine-Tuning MIA in advanced machine learning models, they overlook a critical aspect: both data from the pre-training dataset and fine-tuning dataset are technically a member of the target model. Therefore, to identify a member is actually from the more privacy-concerned fine-tuning dataset, besides identifying the target data as a member, which is the focus of the existing studies, it is also necessary to reject it from the pre-training dataset. One may suggest that this can be easily achieved by conducting MIA again on the original model to reject the target’s member identity. However, this approach has several limitations. Firstly, the original pre-trained model is not available under the black-box setting, but the practical gray box setting is still suitable for real-world simulation and has such access. More importantly, various research has suggested that the fine-tuning process affects the overall MIA vulnerability of the model [41, 51, 52], making the pre-trained model less suitable for calibration. Future research may explore new attack techniques that can effectively separate fine-tuning members from pre-training members, enhancing the robustness of fine-tuning MIA.
Lack of Research in Privacy risks of LMMs Our survey reveals that research on MIA against multimodal models, particularly models beyond VLMs, remains highly limited and lacks advanced attack algorithms. As LMMs continue to expand rapidly across various applications, future research may focus on developing new MIA techniques suitable for the evaluation of LMMs.
7 Conclusion
With the rapid advancement of technology, large-scale machine learning models, including Large Language Models (LLMs) and multimodal models, are increasingly being deployed across various domains. In this work, we covered most existing studies about MIA against LLMs and multimodal models, systematically reviewing each study, categorizing different MIA approaches, and analyzing their strengths and limitations. Furthermore, given the widespread adoption of fine-tuning techniques in model deployment, we also examined current fine-tuning MIA against LLMs and multimodal models. Based on the research gaps identified in the survey, we proposed several directions for future research, aiming to improve the practicality, effectiveness, and robustness of MIAs in real-world scenarios. We hope that this work serves as a valuable resource for researchers and developers working in machine learning security and privacy, fostering further advancements in the field.